OpenStudy (anonymous):

The base of a square pyramid has a side length of 9 inches. The height of the pyramid is 21 inches. What is the volume of the pyramid?

OpenStudy (anonymous):

@paki would I do the same v=bh/3 formula for this one?

OpenStudy (anonymous):

volume of pyramid is (1/3)*area of square base*height

OpenStudy (paki):

ypu

OpenStudy (paki):

solve here.... so we may help ....

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

9*21 = 189 / 3 = 63?

OpenStudy (anonymous):

I don't think that's right

OpenStudy (anonymous):

length of side is given first determine area of square base

OpenStudy (paki):

first find base length.... here it is 9inches... so 9*9=81.... ok... and we have hieght of 21 inches..... now solve....

OpenStudy (anonymous):

k

OpenStudy (paki):

use the same formula... ok... solve here....

OpenStudy (anonymous):

567?

OpenStudy (paki):

how... steps...?

OpenStudy (anonymous):

81*21 = 1701 / 3 = 567

OpenStudy (paki):

yup... excellent... that's the answer...

OpenStudy (anonymous):

Thanks so much!

OpenStudy (paki):

pleasure....